DIFFERENTIAL GEOMETRY/PDE SEMINAR

Tuesday, September 12 Padelford C-401 2:00 PM

Homogeneous hypersurfaces

Michael Eastwood (University of Adelaide)

A homogeneous hypersurface in affine space is one that everywhere looks the same from the affine point of view. I shall talk about their classification in 2, 3, and 4 dimensions. This is joint work with Vladimir Ezhov. The methods are elementary and I shall start from scratch.